Solve the equation in the interval [0°, 360°).
4 sin^2θ = 3

csc θ = 1 + cot θ

3 sin^2θ - sin θ - 4 = 0

2 cos^3θ = cos θ
Answer #1

1.
4\sin^2\theta \:=\:3
3 since 4
\sin \left(\theta\right)=\frac{\sqrt{3}}{2},\:\sin \left(\theta\right)=-\frac{\sqrt{3}}{2}
\\\sin \left(\theta \right)=\frac{\sqrt{3}}{2}\quad :\quad \theta =\frac{\pi }{3}, \frac{2\pi }{3}\\ \sin \left(\theta \right)=-\frac{\sqrt{3}}{2}\quad :\quad \theta =\frac{4\pi }{3}, \frac{5\pi }{3}
{\color{Blue} \theta =\frac{\pi }{3}, \frac{2\pi }{3}, \frac{4\pi }{3}, \frac{5\pi }{3}}
or,
{\color{Blue} \theta =60^{\circ\:},120^{\circ \:}, 240^{\circ \:}, 300^{\circ \:}}

2.
\csc \theta \:=\:1+\cot \theta
\\\frac{1}{\sin \theta} = 1 + \frac{\cos \theta }{\sin \theta}\\
Dinomiator can't be Zero for real solution, \sin \theta \neq 0
1 = \cos \theta + \sin \theta
\\\mathrm{\frac{1}{\sqrt 2} = \frac{1}{\sqrt2} \cos \theta} + \frac{1}{\sqrt 2} *\sin \theta \\ \cos \frac{\pi}{4} = \sin\frac{\pi}{4}* \cos \theta + \cos \frac{\pi}{4}*\sin \theta \\ \cos \frac{\pi}{4} = \cos\left ( \theta - \frac{\pi}{4} \right )
\frac{\pi}{4} = \theta - \frac{\pi}{4}
\\{\color{Blue} \theta = \frac{\pi}{2}}\\ {\color{Blue} \theta = 90\degree}

3.
\\3\:\sin^2\theta \:-\:\sin\:\theta \:-\:4\:=\:0\\ \left (\sin \left(\theta\right) -\frac{4}{3} \right )\:\left (\sin \left(\theta\right) +1 \right ) = 0\\ \sin \left(\theta\right)=\frac{4}{3},\:\sin \left(\theta \right)=-1
\sin \left(\theta \right)=\frac{4}{3}\quad > 1 :\quad \mathrm{No\; solution}
\sin \left(\theta \right)=-1\quad :\quad \theta =\frac{3\pi }{2}
\\ {\color{Blue} \theta = \frac{3\pi }{2}}\\\\ {\color{Blue} \theta = 270 \degree}

4.
\\\cos ^3\theta =\:\cos \theta \\ \cos^3\theta -\cos \theta = 0\\ \cos \theta(\cos^2 \theta - 1) = 0\\ \cos \theta = 0, \cos \theta = -1, \cos \theta = 1
\cos \left(\theta \right)=0\quad :\quad \theta =\frac{\pi }{2},\frac{3\pi }{2}
\cos \left(\theta \right)=-1\quad :\quad \theta =\pi
\cos \left(\theta \right)=1\quad :\quad \theta = 0

{\color{Blue} \theta = 0,\frac{\pi}{2}, \pi, \frac{3\pi}{2}}
or,
{\color{Blue} \theta =0\degree, 90\degree, 180\degree, 270\degree}